Question 4 (20 Marks)

The table below uses data for the year 2000 provide by Statistics Canada and adjusted to be comparable to U.S. data. All values are in millions.

  1. Fill in the blank entries in the table. Show your work

Country

Adult
Population

Labor
Force

Employed

Unemployed

Unemployment
Rate

Labor-Force
Participation
Rate

Japan

108.12

63.79

3.20

France

25.98

2.45

56.4 %

Germany

69.17

39.75

8.08 %

  1. What are some of the types of unemployment discussed in your learning materials.  What are some of the costs of such unemployment?  Is some types of unemployment more troublesome than others? Why or why not?

c. What policies can governments implement to reduce unemployment as close to the natural rate of unemployment as possible? (20 marks)
Answer #1

Hi,

I have only solved the first question because answering the first part required so much time that I could not answer the other parts. I hope this helps!

Thank you.

The table below uses data for the year 2000 provide by Statistics Canada and adjusted to be comparable to U.S. data. All values are in millions.

Country

Adult
Population

Labor
Force

Employed

Unemployed

Unemployment
Rate

Labor-Force
Participation
Rate

Japan

108.12

66.99

63.79

3.20

4.78%

61.96%

France

46.06

25.98

23.53

2.45

9.43%

56.4 %

Germany

69.17

39.75

36.54

3.21

57.48%

Japan:

Labor force for any country is the sum of the number of employed and unemployed people.

Therefore, labor force for Japan = Employed + Unemployed = 63.79 + 3.20 = 66.99

Unemployment Rate for Japan = (Number of Unemployed)/Labor Force * 100 = 3.20/66.99 * 100 = 4.78%

Labor-Force Participation Rate for Japan = Labor Force / Adult Population * 100 = 66.99 / 108.12 * 100 = 61.96%

France:

Now, in France, if Labor Force = 25.98 and Unemployed = 2.45, then number of employed = 25.98 – 2.45 = 23.53

Unemployment Rate for France = (Number of Unemployed)/Labor Force * 100 = 2.45 /25.98 * 100 = 9.43%

Labor-Force Participation Rate for France = 56.4%.

Now we know that, Labor-Force Participation Rate = Labor Force / Adult Population * 100

Therefore, Adult Population for France = Labor Force / Labor-Force Participation Rate * 100 = 46.06

Germany:

Unemployment Rate in Germany = 8.08%

Unemployment Rate for Germany = (Number of Unemployed)/Labor Force * 100

Therefore, we can derive that, number of Unemployed = (Unemployment rate * Labor Force) / 100 = (8.08*39.75)/100 = 3.21

Therefore, employed = Labor force – Unemployed = 39.75 – 3.21 = 36.54

Labor-Force Participation Rate for Germany = Labor Force / Adult Population * 100 =39.75 / 69.17 * 100 = 57.48%